OpenStudy (anonymous):

Four times the sum of a number and 15 is at least 120. Let x represent the number. Find all possible values for x.

OpenStudy (kinggeorge):

\[4(x+15) \geq 120 \Leftrightarrow 4x+60 \geq120\]If we solve for x, we find that \[x \geq15\]

OpenStudy (anonymous):

You drop a rock off a bridge. The rock’s height, h (in feet above the water), after t seconds is modeled by . What is the height of the rock after 2 seconds? A. –64 feet B. 605 feet C. 509 feet D. 477 feet

OpenStudy (anonymous):

Solve. + 9 = 4 A. 65 B. –25 C. 5 D. 20

OpenStudy (kinggeorge):

I'm not here to do your homework for you. I'm here to help you understand how to do your homework. Also, you haven't given me enough information to work off of in your second two questions.

OpenStudy (anonymous):

the seccond question is x/5 +9=4

OpenStudy (kinggeorge):

First, you would have to subtract 9 from both sides to get an isolated x term on one side. Then, since you have x/5, multiply both sides by 5.

OpenStudy (kinggeorge):

\[{x \over 5}+9=4\]\[{x \over 5}=4-9=-5\]\[x=(-5)(5)=-25\]